### v1

@@ 5.2.1 Model performance plateau
Using the ABEA training dataset, the GRACE model was fine-tuned based on the highest average validation scores in the HPO configuration. Nevertheless, the training and validation results for ABEA consistently showed a clear upper boundary in terms of its performance on the joint ATE and AEC tasks. The 10-fold cross-validation yielded an F1-score of 48.0% for the joint task, while the glsATE task alone achieved 69.1%. By comparison, Luo et al. [2] reported 58.3% for the joint tasks and 75.7% for the glsATE-only task on the ABSA Twitter dataset from Mitchell et al. [35]. On the surface, these results, i.e., 10.3 and 5.8 point drops in comparative performance, can be intuitively attributed to the increased complexities of both the new dataset and the ABEA task. The training dataset consisted of more diverse word classes for aspect terms and was less domain-bound than conventional ABSA training datasets (cf. Figure 7). In addition, the identification of emotions rather than sentiment polarities introduced more granularity. Although these changes increase model learning difficulty and can reasonably be expected to cause a drop in performance relative to the ABSA baseline, a closer inspection of the training results provides deeper insights into the underlying causes for the model’s evident performance plateau.
Following the HPO, the cumulative results revealed that the GRACE model was not learning generalizable patterns for the two ABEA tasks, with training performance logs indicating a strong tendency towards overfitting on the training data. This tendency was apparent from a substantial difference between training and validation performance and could be traced back to the original hyperparameters and training settings reported by Luo et al. [2] as optimal for the ATE and ASC tasks, which also suffered from significant overfitting (cf. Figure 8). To reduce overfitting, several configurations were tested during HPO. Most effective among these were the increased dropout rates in BERT ’s hidden layers and the attention heads. At a dropout rate of 0.3, i.e., any given node connection in its network layers with 30% probability is dropped, the model’s training performance plateaued at F1-scores of 73% for the ATE task only and 61% for the joint ATE and AEC tasks, respectively. While this level of dropout successfully prevented overfitting, it did not lead to a more robust performance. Instead, validation metrics significantly dropped, indicating signs of underfitting, as reflected in low performance across both training and validation data.
@@
In light of this evident performance plateau, a critical examination of the training dataset itself is essential Since measures to ensure label consistency were taken, the weakness is therefore likely connected to data quantity [77]. Mosbach et al. [78] argue that BERT fine-tuning instability is not caused by small training datasets themselves, but rather by the reduction of training iterations which often accompany the use of smaller datasets. However, in this study, several settings for increased training iterations did not yield significant performance improvements. A small dataset like the 2,621-instance ABEA dataset likely hinders model learning for complex tasks by failing to provide a sufficiently diverse range of examples, resulting in narrow representations that do not capture the variability needed for real-world application. This can lead to overfitting, where the model memorizes specific patterns in the training dataset, or underfitting when regularization is applied due to a lack of representative data.
@@ 5.2.2 Hyperparameter settings
Various GHL configurations were tested, including its application at different training steps, adjustments to the number of bins (used to group and assign weights to training examples based on their difficulty), and modifications to momentum values. While increasing the number of bins from 24 to 32 yielded the highest average performance by allowing finer differentiation of training examples based on difficulty, none of the GHL adaptations resulted in a significant improvement over previously achieved performances. As a result, the original GHL settings were retained. However, the underlying assumption of the GHL mechanism should be critically evaluated in and of itself. GHL assumes that the most frequent labels are also the easiest to learn. However, this may not apply to emotion classification, where the class distribution does not necessarily correlate with learning difficulty. In the new ABEA dataset, it is not at all self-evident that the most frequent classes, Happiness and None, are easier to learn compared to the least frequent classes Fear and Sadness. In contrast, for the original ABSA dataset [35], GHL is far more applicable since 70% of all examples belong to the Neutral sentiment class.
@@ 6 Conclusion
This paper addresses existing research gaps in the field of Aspect-based Emotion Analysis by creating a consistent training dataset (RQ1) as well as adapting and fine-tuning the GRACE model [2] for the joint extraction of the sub-tasks ATE and AEC (RQ2).
@@
To repurpose the ABSA model GRACE [2] for ABEA, extensive HPO was conducted. The most impactful hyperparameter changes on model performance included reductions in batch size, a warm-up method adjustment, a reduction in the number of shared layers between the ATE and AEC branches, an increase of AEC branch layers, extending training epochs, and omitting VAT in the second training step. A more complex classification layer negatively impacted model performance. The fine-tuned ABEA model plateaued at F1-scores of 70.1% for ATE and 46.9% for joint ATE and AEC.
The lower performance of EmoGRACE for ABEA compared to the reported performance of GRACE for ABSA can be largely attributed to three factors. First, the shift from sentiment polarity to more complex emotion classification. Second, the increased variance in types of words for aspect terms, extending beyond just nouns. Third, the size of the training dataset proved insufficient to allow for generalizable model learning at the high level of task complexity, indicating that data augmentation may be a necessary pre-requisite for continued ABEA research. Future work may consider the use of generative foundation models, such as Generative Pre-Trained Transformer (GPT) models [79], which have shown strong capabilities in handling a variety of NLP tasks and may even outperform fine-tuned task-specific BERT models in certain applications [80].
The effectiveness of the presented ABEA method in capturing public emotions should be evaluated through a targeted, domain-specific case study, such as analyzing posts during a specific natural disaster. The model could be applied to the social media-based analysis of other highly emotional social topics, such as public health crises [81], climate change [82], urban events [59] or political discourses [83], providing insights into collective emotions and their targets.
@@ Acknowledgements
Table 7: Training datasets found in relation to the search terms “aspect-based emotion analysis”, “aspect-based emotion dataset”, or “fine-grained emotion dataset”.
@@
[11] Yes Adidas Instagram account (images and comments) 4,900 Multimodal dataset
[43] No Twitter 5,000 Dataset and model serve the understanding of financial/stock market risk and opportunity
[40] No SemEval Restaurant dataset [31] 3,850 Multilingual, mostly Indonesian language; After IAA the authors removed emotions ‘anger’, ‘fear’ and ‘disgust’ due to small cohorts
[26] No Indonesian restaurant review platform ‘PergiKuliner’ 14,103 Multilingual (English, Dutch, German, French); designed for a business platform
[41] No Bol.com, Trustpilot, TripAdvisor, and proprietary data from project partners 1,000+ Hungarian language
[42] No Hungarian speeches from 2014 - 2018 1,008 Aspects are not labeled per text, but considered as overall categories for which emotion scores are identified; emotions are based on COVID-19 death rates
[84] No Twitter unclear Aspects are not labeled per text, but considered as overall categories for which emotion scores are identified; emotions are based on COVID-19 death rates
@@
Figure 8: Left: Validation metrics per epoch, Right: Training metrics per epoch using the original ABSA dataset with the original hyperparameters and training settings as reported by Luo et al. [2]

### v2

@@ 5.2.1 Model performance plateau
Using the ABEA training dataset, the GRACE model was fine-tuned based on the highest average validation scores in the HPO configuration. Nevertheless, the training and validation results for ABEA consistently showed a clear upper boundary in terms of its performance on the joint ATE and AEC tasks. The 10-fold cross-validation yielded an F1-score of 48.0% for the joint task, while the glsATE task alone achieved 69.1%. By comparison, Luo.2020 reported 58.3% for the joint tasks and 75.7% for the glsATE-only task on the ABSA Twitter dataset from Mitchell.2013. On the surface, these results, i.e., 10.3 and 5.8 point drops in comparative performance, can be intuitively attributed to the increased complexities of both the new dataset and the ABEA task. The training dataset consisted of more diverse word classes for aspect terms and was less domain-bound than conventional ABSA training datasets (cf. Figure 7). In addition, the identification of emotions rather than sentiment polarities introduced more granularity. Although these changes increase model learning difficulty and can reasonably be expected to cause a drop in performance relative to the ABSA baseline, a closer inspection of the training results provides deeper insights into the underlying causes for the model’s evident performance plateau.
Following the HPO, the cumulative results revealed that the GRACE model was not learning generalizable patterns for the two ABEA tasks, with training performance logs indicating a strong tendency towards overfitting on the training data. This tendency was apparent from a substantial difference between training and validation performance and could be traced back to the original hyperparameters and training settings reported by Luo.2020 as optimal for the ATE and ASC tasks, which also suffered from significant overfitting (cf. Figure 8). To reduce overfitting, several configurations were tested during HPO. Most effective among these were the increased dropout rates in BERT ’s hidden layers and the attention heads. At a dropout rate of 0.3, i.e., any given node connection in its network layers with 30% probability is dropped, the model’s training performance plateaued at F1-scores of 73% for the ATE task only and 61% for the joint ATE and AEC tasks, respectively. While this level of dropout successfully prevented overfitting, it did not lead to a more robust performance. Instead, validation metrics significantly dropped, indicating signs of underfitting, as reflected in low performance across both training and validation data.
@@
In light of this evident performance plateau, a critical examination of the training dataset itself is essential Since measures to ensure label consistency were taken, the weakness is therefore likely connected to data quantity [Ezen-Can.2020]. Mosbach.2021 argue that BERT fine-tuning instability is not caused by small training datasets themselves, but rather by the reduction of training iterations which often accompany the use of smaller datasets. However, in this study, several settings for increased training iterations did not yield significant performance improvements. A small dataset like the 2,621-instance ABEA dataset likely hinders model learning for complex tasks by failing to provide a sufficiently diverse range of examples, resulting in narrow representations that do not capture the variability needed for real-world application. This can lead to overfitting, where the model memorizes specific patterns in the training dataset, or underfitting when regularization is applied due to a lack of representative data.
@@ 5.2.2 Hyperparameter settings
Various GHL configurations were tested, including its application at different training steps, adjustments to the number of bins (used to group and assign weights to training examples based on their difficulty), and modifications to momentum values. While increasing the number of bins from 24 to 32 yielded the highest average performance by allowing finer differentiation of training examples based on difficulty, none of the GHL adaptations resulted in a significant improvement over previously achieved performances. As a result, the original GHL settings were retained. However, the underlying assumption of the GHL mechanism should be critically evaluated in and of itself. GHL assumes that the most frequent labels are also the easiest to learn. However, this may not apply to emotion classification, where the class distribution does not necessarily correlate with learning difficulty. In the new ABEA dataset, it is not at all self-evident that the most frequent classes, Happiness and None, are easier to learn compared to the least frequent classes Fear and Sadness. In contrast, for the original ABSA dataset [Mitchell.2013], GHL is far more applicable since 70% of all examples belong to the Neutral sentiment class.
@@ 6 Conclusion
This paper addresses existing research gaps in the field of Aspect-based Emotion Analysis by creating a consistent training dataset (RQ1) as well as adapting and fine-tuning the GRACE model [Luo.2020] for the joint extraction of the sub-tasks ATE and AEC (RQ2).
@@
To repurpose the ABSA model GRACE [Luo.2020] for ABEA, extensive HPO was conducted. The most impactful hyperparameter changes on model performance included reductions in batch size, a warm-up method adjustment, a reduction in the number of shared layers between the ATE and AEC branches, an increase of AEC branch layers, extending training epochs, and omitting VAT in the second training step. A more complex classification layer negatively impacted model performance. The fine-tuned ABEA model plateaued at F1-scores of 70.1% for ATE and 46.9% for joint ATE and AEC.
The lower performance of EmoGRACE for ABEA compared to the reported performance of GRACE for ABSA can be largely attributed to three factors. First, the shift from sentiment polarity to more complex emotion classification. Second, the increased variance in types of words for aspect terms, extending beyond just nouns. Third, the size of the training dataset proved insufficient to allow for generalizable model learning at the high level of task complexity, indicating that data augmentation may be a necessary pre-requisite for continued ABEA research. Future work may consider the use of generative foundation models, such as Generative Pre-Trained Transformer (GPT) models [Radford.2018], which have shown strong capabilities in handling a variety of NLP tasks and may even outperform fine-tuned task-specific BERT models in certain applications [Zhong.2023].
The effectiveness of the presented ABEA method in capturing public emotions should be evaluated through a targeted, domain-specific case study, such as analyzing posts during a specific natural disaster. The model could be applied to the social media-based analysis of other highly emotional social topics, such as public health crises [Kogan.2021], climate change [Li.2023a], urban events [Honzak.2024] or political discourses [Kovacs.2021], providing insights into collective emotions.
@@ Acknowledgements
Table 7: Training datasets found in relation to the search terms “aspect-based emotion analysis”, “aspect-based emotion dataset”, or “fine-grained emotion dataset”. The dataset used in this paper is featured in the last row.
@@
[DeBruyne.2022] Yes Adidas Instagram account (images and comments) 4,900 Multimodal dataset
[Garcia-Mendez.2023] No Twitter 5,000 Dataset and model serve the understanding of financial/stock market risk and opportunity
[Maitama.2021] No SemEval Restaurant dataset [Pontiki.2014] 3,850 Multilingual, mostly Indonesian language; After IAA the authors removed emotions ‘anger’, ‘fear’ and ‘disgust’ due to small cohorts
[Suciati.2020] No Indonesian restaurant review platform ‘PergiKuliner’ 14,103 Multilingual (English, Dutch, German, French); designed for a business platform
[DeGeyndt.2022] No Bol.com, Trustpilot, TripAdvisor, and proprietary data from project partners 1,000+ Hungarian language
[Uveges.2022] No Hungarian speeches from 2014 - 2018 1,008 Aspects are not labeled per text, but considered as overall categories for which emotion scores are identified; emotions are based on COVID-19 death rates
[DeviT..2021] No Twitter unclear Aspects are not labeled per text, but considered as overall categories for which emotion scores are identified; emotions are based on COVID-19 death rates
Yes Twitter 2,621 English dataset with complex aspect term annotation, four emotions and a None class.
@@
Figure 8: Left: Validation metrics per epoch, Right: Training metrics per epoch using the original ABSA dataset with the original hyperparameters and training settings as reported by Luo.2020